### v2

# Quadrupole-hexadecapole correlations in neutron-rich samarium and gadolinium isotopes

Abstract
We present an extensive study of quadrupole-hexadecapole correlation effects in even-even Sm and Gd isotopes with neutron number $N=88-106$. The calculations are performed in the framework of the Gogny energy density functional (EDF) with the D1S parametrization and the $sdg$ interacting boson model (IBM). The quadrupole-hexadecapole constrained self-consistent mean-field potential energy surface is mapped onto the expectation value of the $sdg$-boson Hamiltonian. This procedure determines the parameters of the $sdg$-IBM Hamiltonian microscopically. Calculated excitation energies and transition strengths are compared to the ones obtained with a simpler $sd$-IBM, as well as with the experimental data. The Gogny-EDF mapped $sdg$-IBM reproduces spectroscopic properties of the studied nuclei as reasonably as in the case of the previous $sdg$-boson mapping calculations that were based on the relativistic EDF, indicating that the axial quadrupole-hexadecapole method is sound regardless of whether relativistic or nonrelativistic EDF is employed. The mapped $sdg$-IBM improves some of the results in lighter Sm and Gd isotopes compared to the mapped $sd$-IBM, implying the existence of significant hexadecapole correlations in those nuclei. For those nuclei with $N\geq 94$, hexadecapole effects are minor, and the only significant difference between the two boson models can be found in the description of $E0$ monopole transitions.
I Introduction
The importance of studying deformations in nuclei and their effects on various nuclear properties, e.g. excitation energies and transition strengths, has been recognized for decades [1, 2]. The most commonly present and extensively studied nuclear deformation is of quadrupole type, and its effects on the positive-parity states are well known. The higher-order correlations which affect the positive-parity states, the hexadecapole correlations, have been significantly less studied, due to their effects often being overshadowed by large quadrupole correlations. Nevertheless, the presence of hexadecapole correlations has been found in a wide range of nuclei [3, 4], and the main effect of such correlations is the appearance of the $4^{+}$ band in the low-lying excitation energy spectra of some even-even nuclei, with an enhanced $B(E4)$ transition strength from the $4^{+}$ bandhead state to the $0^{+}$ ground state. Hexadecapole correlations have also been found to be important in describing certain nuclear reactions and decays [4, 5, 6].
The interacting boson model (IBM) [7] is one of the most commonly employed models for studying how nuclear deformations affect low-lying spectra [8]. The main assumption of the model is that a nucleus can be approximated as a system consisting of a core, represented by the closest doubly-magic nucleus, and the valence nucleons, which form collectives pairs (or bosons). Interactions among valence bosons give rise to the low-lying spectra and transitions. In the simplest version of the model, the $sd$-IBM-1, the nucleons primarily couple into $s$ $(J^{\pi}=0^{+})$ and $d$ $(J^{\pi}=2^{+})$ bosons, and the neutron and proton bosons are assumed to be identical [7]. One particular extension of the simple $sd$-IBM is the inclusion of $g$ bosons $(J^{\pi}=4^{+})$, allowing for the study of hexadecapole collectivity in nuclei. The importance of $g$ bosons in describing low-lying collective states has been extensively studied [8, 9, 10, 11, 12, 13, 14, 15]. In recent years, a method has been developed, which is to derive the parameters of the IBM Hamiltonian from self-consistent mean-field (SCMF) calculations based on a microscopic energy density functional (EDF) [16]. The method provides low-energy collective energy spectra and reduced transition probabilities, and has been successfully applied to study quadrupole [16, 17, 18, 19], octupole [20, 21], and, recently, hexadecapole correlations [22, 23]. In this paper, we present an extension of Refs. [22, 23] by studying the low-energy quadrupole and hexadecapole collective states of even-even rare-earth isotopes 150-168Sm and 152-170Gd, ranging from transitional to highly deformed nuclei, based on the Gogny force [24]. The present calculation employs parametrization D1S [25] of the Gogny interaction, which has already been extensively employed [26, 27]. The combination of the Gogny EDFs with the IBM through the mapping procedure has been considered and shown to be valid to describe a number of nuclear structure phenomena, including coexistence of quadrupole shapes (e.g., Refs. [28, 29]) and octupole correlations (e.g., Refs. [30, 31, 32, 33]).
Recently, Gogny Hartree-Fock-Bogoliubov (HFB) calculations, with constraints set on the quadrupole and hexadecapole deformation parameters, were performed to compute potential energy surfaces (PESs) in rare-earth nuclei [34]. In this study, we carry out the mapping procedure by using the Gogny-D1S PESs in order to examine the effects of the quadrupole-hexadecapole coupling on the low-lying spectra of the considered nuclei. Since, in previous studies [22, 23], calculations were performed within the framework of the constrained relativistic mean field (MDC-RMF) model [35, 36], with the relativistic density-dependent point-coupling (DD-PC1) functional [37, 38], this work will allow us to study how the quadrupole-hexadecapole mapping method depends on the choice of the EDF. On the other hand, we analyze positive-parity bands in heavy Sm and Gd isotopes as well as the effects of hexadecapole collectivity in heavy neutron-rich nuclei, far from the valley of stability. The results of the calculations are compared with the ones obtained with a simpler $sd$-IBM, as well as with the available experimental data [39].
The paper is organized as follows. The theoretical framework, i.e., the Gogny-D1S HFB approach and the $sd$-IBM and $sdg$-IBM, is outlined in Sec. II. Potential energy surfaces computed with both SCMF and IBM calculations are presented in Sec. III. Spectroscopic properties, including the excitation spectra of low-lying states, electric quadrupole, hexadecapole, and monopole transition properties, are discussed in Sec. IV. Section V is devoted to the concluding remarks and work perspectives.
II Model description
In order to obtain the SCMF PES for the studied nuclei, the HFB equations have been solved with constraints on the axial quadrupole $\hat{Q}_{20}$ and hexadecapole $\hat{Q}_{40}$ moment operators [34]. The axial quadrupole $\beta_{20}\equiv\beta_{2}$ and hexadecapole $\beta_{40}\equiv\beta_{4}$ deformation parameters are related to the expectation values of the multipole moment operators as [34]
$\beta_{\lambda}=\frac{\sqrt{4\pi(2\lambda+1)}}{3R^{\lambda}A}\left\langle\hat{Q}_{\lambda 0}\right\rangle,$ (1)
where $\lambda=2$ or 4, $R=1.2A^{1/3}$ fm and $A$ is the nucleon number. Details on the Gogny-D1S HFB calculations can be found in Refs. [25, 40, 41].
In order to calculate low-lying excitation energies and transition strengths, we have employed the $sdg$-IBM-1, referred to simply as $sdg$-IBM. The Hamiltonian is built in the same way as in Ref. [23]:
$\hat{H}=\epsilon_{d}\hat{n}_{d}+\epsilon_{g}\hat{n}_{g}+\kappa\hat{Q}^{(2)}\cdot\hat{Q}^{(2)}+\kappa(1-\chi^{2})\hat{Q}^{(4)}\cdot\hat{Q}^{(4)},$ (2)
with $\hat{n}_{d}=d^{\dagger}\cdot\tilde{d}$ and $\hat{n}_{g}=g^{\dagger}\cdot\tilde{g}$ being the $d$- and $g$-boson number operators, respectively,
$\hat{Q}^{(2)}=$ $(s^{\dagger}\times\tilde{d}+d^{\dagger}\times s)+\chi\Big{[}\frac{11\sqrt{10}}{28}(d^{\dagger}\times\tilde{d})^{(2)}$ (3)
$-\frac{9}{7}\sigma(d^{\dagger}\times\tilde{g}+g^{\dagger}\times\tilde{g})^{(2)}+\frac{3\sqrt{55}}{14}(g^{\dagger}\times\tilde{g})^{(2)}\Big{]}$
being the quadrupole operator, and
$\hat{Q}^{(4)}=s^{\dagger}\times\tilde{g}+g^{\dagger}\times s$ (4)
[…]
The parameters $\epsilon_{d}$, $\epsilon_{g}$, $\kappa$, $\chi$ and $\sigma$ are determined by the mapping procedure [16]. The $sdg$-IBM is connected to the geometric model through a coherent state [43], |$\ket{\phi}\sim(1+\tilde{\beta}_{2}d^{\dagger}_{0}+\tilde{\beta}_{4}g^{\dagger}_{0})^{N_{B}}\ket{0}$, with $N_{B}$ representing the number of bosons, determined by the nearest doubly magic nucleus, and |$\ket{0}$ representing the boson vacuum. Note that $\tilde{\beta}_{2}$ and $\tilde{\beta}_{4}$ represent the axial bosonic quadrupole and hexadecapole deformation parameters, respectively. For Gd and Sm isotopes with $N<106$, the core is represented by the doubly magic nucleus 132Sn. In the case of 168Sm and 170Gd, the corresponding core is 176Sn since the neutrons occupy the upper half of the neutron major shell $N=82-126$.
The parameters of the $sdg$-IBM are fitted so that the PES of the IBM, $E_{\textnormal{IBM}}(\tilde{\beta}_{2},\tilde{\beta}_{4})=\bra{\phi}\hat{H}\ket{\phi}/\braket{\phi}{\phi}$, should be approximately equal to the Gogny-HFB PES in the vicinity of the minimum:
$E_{\textnormal{SCMF}}(\beta_{2},\beta_{4})\approx E_{\textnormal{IBM}}(\tilde{\beta}_{2},\tilde{\beta}_{4})\;.$ (5)
The relation between axial bosonic and fermionic parameters is assumed to be linear, $\tilde{\beta}_{2}=C_{2}\beta_{2}$, $\tilde{\beta}_{4}=C_{4}\beta_{4}$ as in previous studies with the mapping procedure [16, 21, 22, 23]. This leaves us with 7 parameters which need to be fitted to reproduce the SCMF PES.
For the $sd$-IBM calculations, we have employed the following Hamiltonian [7]:
$\hat{H}_{sd}=\epsilon_{d}\hat{n}_{d}+\kappa\hat{Q}^{(2)}\cdot\hat{Q}^{(2)},$ (6)
where the $sd$-IBM quadrupole operator reads
$\hat{Q}^{(2)}=s^{\dagger}\tilde{d}+d^{\dagger}s+\chi\left(d^{\dagger}\times\tilde{d}\right)^{(2)}.$ (7)
The parameters $\epsilon_{d}$, $\kappa$, $\chi$ in the $sd$-IBM Hamiltonian are fixed by a 1D mapping of the Gogny-D1S PES along the $\beta_{4}=0$ line:
$E_{\textnormal{SCMF}}(\beta_{2},\beta_{4}=0)=E_{sd-\textnormal{IBM}}(\tilde{\beta}_{2}),$ (8)
and the relation between the bosonic and fermionic deformation parameters is assumed to be linear $\tilde{\beta}_{2}=C^{sd}_{2}\beta_{2}$.
In order to calculate transition strengths with both the $sdg$-IBM and $sd$-IBM, the transition operators are constructed in the same way as in Refs. [22, 23]. The quadrupole transition operator is of the form:
$\hat{T}(E2)_{sdg/sd}=e_{2}^{sdg/sd}\hat{Q}^{(2)}_{sdg/sd},$ (9)
with $\hat{Q}^{(2)}_{sdg/sd}$ representing the quadrupole operator of the $sdg$ (3) or $sd$ (7) boson model. The hexadecapole operator is chosen, in the case of $sdg$-IBM, to be
$\hat{T}^{(E4)}_{sdg}=e_{4}^{sdg}\left[s^{\dagger}\tilde{g}+g^{\dagger}s+(d^{\dagger}\times\tilde{d})^{(4)}\right],$ (10)
whereas for $sd$-IBM it is constructed as
$\hat{T}^{(E4)}_{sd}=e_{4}^{sd}(d^{\dagger}\times\tilde{d})^{(4)}.$ (11)
The quadrupole and hexadecapole effective charges $e_{2}^{sdg/sd}$ and $e_{4}^{sdg/sd}$ are fitted in order to reproduce the measured transition strengths $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ and $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$, from the first $2^{+}$ and $4^{+}$ states to the ground state $0^{+}$, respectively. The monopole transition operator [44] reads
$\hat{T}^{(E0)}_{sdg}=(e_{n}N+e_{p}Z)\left(\eta\frac{\hat{n}_{d}}{N_{B}}+\gamma\frac{\hat{n}_{g}}{N_{B}}\right),$ (12)
for the $sdg$-IBM, and
$\hat{T}^{(E0)}_{sd}=(e_{n}N+e_{p}Z)\eta\frac{\hat{n}_{d}}{N_{B}},$ (13)
for the $sd$-IBM.
Following Ref. [44], we set neutron and proton effective charges to the values $e_{p}=e$ and $e_{n}=0.5e$, while the values of parameters $\eta$ and $\gamma$ are taken to be $\eta=\gamma=0.75$ fm2. Experimental data for energies and transition strengths have been taken from the NNDC database [39].
Figure 1: Axially symmetric quadrupole ($\beta_{20}$) and hexadecapole ($\beta_{40}$) constrained potential energy surfaces for 150-166Sm. The energy difference between neighbouring contours is 0.1 MeV. The absolute minimum is indicated by an open triangle. Results have been obtained with the Gogny-D1S energy density functional.
Figure 2: The same as in Fig. 1, but for 152-168Gd.
Figure 3: The same as in Fig. 1, but for the mapped $sdg$-IBM potential energy surfaces of 150-166Sm.
Figure 4: The same as in Fig. 1, but for the mapped $sdg$-IBM potential energy surfaces of 152-168Gd.
III Mapping the SCMF results onto the IBM space
Figures 1 and 2 show the Gogny-D1S PESs obtained for even-even Sm and Gd isotopes with neutron numbers within the range $N=88-106$, up to an energy of 2.7 MeV above the HFB ground state. The PESs for the $N=106$ isotones are not shown due to their similarity to those of the $N=104$ ones. As can be seen from the figures, the quadrupole and hexadecapole deformations associated with the absolute minima of the PESs start from the values $\beta_{2}^{\textnormal{min}}=0.2$ and $\beta_{4}^{\textnormal{min}}=0.08$, respectively. They increase with neutron number until $N=94$, after which hexadecapole deformations start decreasing, dropping to $\beta_{4}^{\textnormal{min}}=0.04$ at $N=106$. Similar to previous (mapped) IBM calculations, based on the relativistic DD-PC1 functional [22, 23], Sm isotopes exhibit larger HFB ground state $\beta_{4}$ values as compared to Gd isotopes. On the other hand, the ground state $\beta_{2}$ deformations increase up to $N=94$. For larger neutron numbers, those quadrupole deformations decrease, reaching the value $\beta_{2}^{\textnormal{min}}=0.32$ for $N=106$ isotopes. An important thing to observe is that the PESs appear to be ”tilted” with respect to $\beta_{20}$ and $\beta_{40}$ axes, indicating a strong coupling between the axial quadrupole and hexadecapole degrees of freedom, which has already been discussed in Ref. [34]. Note, that the results obtained for $N=88-96$ isotopes agree well with the ones obtained with the DD-PC1 EDF [22, 23].
The corresponding (mapped) $sdg$-IBM PESs are depicted in Figs. 3 and 4. The mapping procedure reproduces basic features of the fermionic PESs such as the position of the absolute minima, the oblate saddle points for $N=88$ nuclei as well as the overall shape of the surface. The strong quadrupole-hexadecapole coupling is well reproduced in $N\leq 100$ isotopes, whereas in $N>100$ isotopes, the coupling is predicted to be less strong. Note that the IBM PESs are significantly flatter than the HFB ones, due to the restricted boson (valence) space as compared to the Gogny-HFB model space. This also explains why the mapping procedure is not able to reproduce the “elongated” shapes of the PESs for $N\geq 100$ nuclei. The 1D $sd$-IBM mapping has also been carried out in all the studied isotopes to reproduce some of the properties of the $E_{\textnormal{SCMF}}(\beta_{2},\beta_{4}=0)$ curve such as the position of the absolute minimum, the position and energy of the oblate saddle point, and the energy of the spherical $\beta_{2}=0$ configuration.
Figure 5: Parameters of the $sdg$ IBM Hamiltonian Eq.(2), as functions of the neutron number $N$.
Figure 6: Parameters of the $sd$ IBM Hamiltonian Eq.(6), as functions of the neutron number $N$.
The $sdg$- and $sd$-IBM parameters are shown in Figs. 5 and 6, as functions of the neutron number $N$. The $d$-boson energy $\epsilon_{d}$ has a very similar behavior in both models. It decreases up to $N=100$, after which it stops decreasing, and even increases at $N=106$, which corresponds to lower $\beta_{2}^{\rm min}$ values. Moreover, the parameters $\kappa$, $\chi$ and $C_{2}$ also display a similar behavior in both models, whereas in the $sdg$-IBM the $\chi$ parameter is characterized by significantly lower absolute values, compared to the $sd$-IBM ones. Regarding those parameters appearing only in the $sdg$-IBM, the $g$-boson energy $\epsilon_{g}$ takes values between $\epsilon_{g}=1.1$ and $\epsilon_{g}=1.3$ MeV. The $C_{4}$ parameter decreases until $N=96$, after which it starts to increase, which corresponds to $\beta_{4}^{\rm min}$ values becoming lower for $N\geq 96$ isotopes. Finally, with increasing neutron number, the parameter $\sigma$, which accounts for the quadrupole-hexadecapole coupling in the $sdg$-IBM, drops from $\sigma=2.8$ to $\sigma=1.0$, which also corresponds to a decrease in $\beta_{4}^{\textnormal{min}}$. This explains why the coupling is predicted to be less strong in $N>100$ isotopes by the $sdg$-IBM. It should also be noted that $\sigma$ decreases faster in Gd isotopes, compared to the Sm ones, which can be explained from the fact that $\beta_{4}^{\textnormal{min}}$ values are smaller in Gd isotopes with the same $N$.
IV Results of the spectroscopic calculations
In this section, we discuss the excitation energies and transition strengths obtained from the diagonalization of the IBM Hamiltonian, using the computer program ARBMODEL [45]. The results of the $sdg$-IBM and $sd$-IBM will be compared to illustrate the effects of $g$ bosons. The results obtained from both models will also be compared with the available experimental data at NNDC [39].
IV.1 Excitation energies
Figure 7: Excitation energies of the yrast band states up to spin $J^{\pi}=14^{+}$ as functions of the neutron number $N$ within the mapped $sdg$-IBM (left column) and $sd$-IBM (right column), represented by solid symbols connected by solid lines. Experimental data are taken from Ref. [39], and are depicted as open symbols connected by dotted lines.
Figure 7 shows the excitation energies corresponding to the yrast band states with even spin $J^{\pi}=2^{+}$ to $14^{+}$. As expected, the $sdg$-IBM improves the description of the excitation energies of higher-spin states with $J^{\pi}=12^{+}$ and $14^{+}$ in transitional $N=88$ nuclei. For Sm nuclei, the $sdg$-IBM gives a better description of the energy levels up to $N=92$ than the $sd$-IBM. This can be explained by the fact that large amounts of $g$-boson contributions are present in the higher-spin states, as indeed, the expectation value of the $g$-boson number operator $\braket{\hat{n}_{g}}$ is greater than or equal to 1 in those isotopes with $N\leq 92$. Note that for Sm isotopes with $N\leq 92$, the $\beta_{4}^{\rm min}$ values are larger compared to those for Gd isotopes with the same $N$. For the $N>92$ isotopes, there is no significant difference between both models concerning the description of the ground state bands. This is not surprising, since as the number of valence nucleons increases for deformed nuclei even the $sd$-IBM is good enough to provide an accurate description of the yrast-band levels. In any case, in general the $sdg$-IBM does provide some significant improvements in the description of the yrast band, compared to the simpler $sd$-IBM.
Figure 8: The same as in Fig. 7, but for the excitation energies of $0_{2}^{+}$, $2_{3}^{+}$, and $4_{3}^{+}$ states.
Figure 8 depicts the excitation energies of the $0_{2}^{+}$, $2_{3}^{+}$ and $4_{3}^{+}$ states, which are associated with the $K^{\pi}=0^{+}$ band in the present study. The $sdg$-IBM predicts a significantly low-lying $4_{3}^{+}$ level for $N=88$ isotopes. This agrees better with the experimental data than in the corresponding $sd$-IBM results. For other isotopes, there is no significant difference between the two models, as they both predict similar energies for all three states. Overall, the $sdg$-IBM does not improve the description of the $0^{+}$ band significantly.
Figure 9: The same as in Fig. 7, but for the excitation energies of $2_{2}^{+}$, $3_{1}^{+}$, and $4_{2}^{+}$ states.
Figure 9 shows the excitation energies of the $2^{+}_{2}$, $3^{+}_{1}$, and $4^{+}_{2}$ states, which can be identified as members of the $\gamma$-vibrational band in the present calculation. Both models predict similar energies for Sm and Gd isotopes. While the calculated energies of the $\gamma$-vibrational band states are in good agreement with the experimental data for the nuclei with $N\leq 90$ and $N\geq 96$, both models underestimate the energies for those with $N=92$ and 94. Overall, the presence of $g$ bosons in the IBM calculations does not seem to have any significant effects on the $\gamma$-vibrational band. The description of the $\gamma$-vibrational band could be improved by including both three-body terms [19] and the $\gamma$ degree of freedom at the HFB level. However, such an extension is beyond the scope of this paper. Work along these lines will be reported in future publications.
IV.2 Transition strengths
IV.2.1 $E2$ transitions
Figure 10: $B(E2)$ transition strengths in the ground state band of the well-deformed $N=90$ (first row) and $N=92$ (second row) isotopes as functions of spin $J$, computed with the mapped $sdg$-IBM (solid curves) and $sd$-IBM (dotted curves). The experimental data, represented by solid circles, are taken from Ref. [39].
Figure 10 shows the $B(E2;J\rightarrow J-2)$ transition strengths in the ground state bands of the well deformed $N=90$ and 92 isotones. These nuclei are specifically considered, since there are experimental data on $E2$ transitions. In Sm isotopes, the $sdg$-IBM significantly improves the description of $B(E2)$ transition strengths for $J^{\pi}\geq 6^{+}$. However, for 154Gd, the $sdg$-IBM predicts $B(E2)$ strengths significantly larger than the $sd$-IBM, which are however still smaller than the measured values. The discrepancy points towards some missing correlations other than the hexadecapole ones, that could also contribute to the $B(E2)$ strengths. Due to the large experimental error bars, in 156Gd it is not possible to conclude whether $sdg$-IBM improves the description of the $E2$ transitions in the yrast band or not. The fact that the $sdg$-IBM predicts larger $B(E2)$ transition strengths than the $sd$-IBM can be attributed to the large values of the parameter $\sigma$ (see Fig. 5). In the case of 152,154Sm and 154Gd the value $\sigma=2.8$ leads to a significant contribution to the $B(E2)$ transition strengths arising from the $\left[d^{\dagger}\times\tilde{g}+g^{\dagger}\times\tilde{d}\right]^{(2)}$ part of the quadrupole $\hat{Q}^{(2)}$ operator of Eq.(3). Moreover, the $\sigma=2.1$ value obtained for 156Gd leads to similar $sdg$-IBM and $sd$-IBM $B(E2)$ strengths. Note that the $\sigma$ values for 154,156Gd are larger than those previously obtained [23] using microscopic input from the relativistic DD-PC1 EDF.
IV.2.2 Hexadecapole transitions
Figure 11: $B(E4;4^{+}_{n}\to 0^{+}_{1})$ ($n=1,2,3,4$) transition strengths as functions of the mass number $A$, computed with the mapped $sdg$-IBM (left column) and $sd$-IBM (right column). Experimental data, indicated by solid circles in the plots, are taken from Refs. [39, 46, 47, 48].
The $B(E4;4_{n}^{+}\rightarrow 0_{1}^{+})$ ($n=1,2,3,4$) reduced transition probabilities obtained in the mapped calculations are plotted in Fig. 11, as functions of the nucleon number $A$ in order to avoid overlapping between the data for Sm and Gd isotopes. The $e_{4}^{sdg/sd}$ effective charges are fitted to the experimental $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$ values [39, 46, 47, 48]. Results for $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$ strengths are plotted in Fig. 11(a) and 11(b), while $B(E4)$ values from non-yrast $4_{n}^{+}$ (with $n=2,3,4$) to the $0^{+}$ ground states are shown in Figs. 11(c)-11(h). The $4_{2}^{+}$ state typically belongs to the $\gamma$-vibrational band, where $g$ bosons do not play a significant role. Thus both models predict weak $E4$ transitions in this case. Large $sdg$-IBM $B(E4)$ values are obtained for the $4_{3}^{+}$ states in 150Sm and 152Gd. The strong $E4$ transition, predicted for the $4_{4}^{+}$ state in 154Gd, agrees well with the experimental identification of this state as a bandhead of the $K=4^{+}$ band [39]. On the other hand, the $sd$-IBM gives negligibly small $B(E4)$ reduced transition probabilities. This indicates a significant difference between the $sdg$-IBM and the simpler $sd$-IBM. Nevertheless, experimental data on $B(E4)$ strengths in rare-earth nuclei are still required to understand whether the $sdg$-IBM describes them better.
IV.2.3 Monopole transitions
Figure 12: $\rho^{2}(E0;0_{i}^{+}\rightarrow 0_{j}^{+})$ values as functions of the neutron number $N$ for Sm and Gd isotopes, computed with the mapped $sdg$-IBM (left column) and $sd$-IBM (right column). Experimental values, plotted as solid circles, are taken from Refs. [39, 49].
Figure 12 shows the monopole strengths $\rho^{2}(E0;0_{i}^{+}\rightarrow 0_{j}^{+})$, with $i=2,3$ and $j=1,2$, as functions of the neutron number $N$. In $N\leq 92$ isotopes, both models yield similar values of monopole strengths. As can be seen from the figure, $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{2}^{+})$ vanishes for $N>92$. The $sdg$-IBM predicts a much sharper decrease (increase) of $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ [$\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$], with increasing neutron number, than the $sd$-IBM.
One can see from Figs. 12(a) and (c) that for those Sm and Gd nuclei with $N\geq 92$, the $sdg$-IBM predicts the monopole strengths $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ and $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ to be more or less similar to each other, whereas in the $sd$-IBM these monopole transition strengths are quite at variance, that is, the $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ are larger than the $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ by a factor of 2 to 3. In addition, the $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ values obtained from the mapped $sdg$-IBM are smaller than those from the $sd$-IBM for neutron-rich nuclei with $N\geq 94$. The reduction of the $0_{2}^{+}\rightarrow 0_{1}^{+}$ $E0$ transition in the case of the $sdg$-IBM has been already been discussed in Ref. [44]. On the other hand, the $sdg$-IBM provides larger $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ values for the $N\geq 94$ nuclei than the $sd$-IBM. These quantitative differences in the predicted $\rho^{2}(E0)$ values between the two sets of the IBM calculations represent the most significant difference between the $sdg$-IBM and $sd$-IBM in the neutron-rich Sm and Gd isotopes with $N\geq 94$. As in the case of the hexadecapole transitions, experimental information about the monopole transitions is limited. It is therefore not possible to tell whether the $sdg$-IBM represents an improvement over the simpler $sd$-IBM in this respect.
V Summary
We have presented an extensive analysis of the quadrupole-hexadecapole collectivity in even-even rare-earth isotopes 150-168Sm and 152-170Gd, and the effects of hexadecapole correlations on low-lying excitation energies and transition strengths. The calculations were performed in the framework of the Gogny D1S EDF and the $sdg$-IBM and $sd$-IBM. The calculated results for the excitation spectra were shown to be in good agreement with the experimental data at the same level of accuracy as our previous mapped $sdg$-IBM studies based on the relativistic EDF [22, 23]. We have thus found that the quadrupole-hexadecapole mapping method does not significantly depend on the choice of the EDF.
The mapped $sdg$-IBM improves the description of the high-spin yrast states in the nuclei with $N=88$, as well as in the Sm isotopes with $N=90$ and 92. In heavier rare-earth nuclei, the contribution of $g$ bosons to the ground state band is insignificant. In the case of the excited $K=0^{+}$ and $\gamma$-vibrational bands, the $sdg$-IBM results did not differ significantly from the $sd$-IBM ones, with only some minor improvements of the $4_{3}^{+}$ states of the $K=0^{+}$ band at $N=88$ in the $sdg$-IBM. In well-deformed $N=90$ and 92 nuclei, the $sdg$-IBM predicts stronger $B(E2)$ transition strengths for the yrast states with spin $J^{\pi}\geq 6^{+}$, which is in good agreement with the experimental data. The effect is more visible in 152,154Sm, for which pronounced hexadecapole mean-field minima were found in the PESs. The choice of the Gogny D1S EDF seems to lead to somewhat stronger yrast band $B(E2)$ values compared to the ones obtained with the relativistic DD-PC1 EDF. The existence of the $K=4^{+}$ band with an enhanced $B(E4;4^{+}\rightarrow 0^{+})$ transition to the ground state is predicted by the $sdg$-IBM in 150Sm and 152,154Gd. In heavier isotopes, such bands are also predicted; however, the bandhead in those nuclei is predicted to be a higher-lying $4_{n\geq 5}^{+}$ state. Regarding monopole strengths, the $sdg$-IBM predicts lower values of $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ and larger values of $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ in neutron-rich isotopes with $N\geq 94$ than the $sd$-IBM. This is in good agreement with previous theoretical calculations, and represents the most pronounced hexadecapole correlation effect on the low-lying spectra in very neutron-rich Sm and Gd isotopes. Due to the lack of experimental data on such transitions, it remains to be seen whether the $sdg$-IBM correctly predicts the behavior of monopole transitions in those isotopes.
Now that the mapping method has been shown to be adequate for describing the quadrupole-hexadecapole collectivity in nuclei, regardless of the choice of the EDF, the method can be extended to include even-odd and odd-odd nuclei. The method should also be applied to heavier rare-earth nuclei up to W and Os nuclei, where more experimental data on $E4$ transitions are available, and to explore possible hexadecapole correlations in other regions of the nuclear chart such as actinides. The method could also be extended to include proton and neutron degrees of freedom (i.e., $sdg$-IBM-2), which become relevant to describe phenomena like mixed-symmetry states and scissor modes.
Acknowledgements.
Most of the work by L.L. was done at the University of Zagreb and was funded within the Tenure Track Pilot Programme of the Croatian Science Foundation and the École Polytechnique Fédérale de Lausanne, and the project TTP-2018-07-3554 Exotic Nuclear Structure and Dynamics, with funds from the Croatian-Swiss Research Programme. This work has also been supported by the Spanish Agencia Estatal de Investigación (AEI) of the Ministry of Science and Innovation (MCIN) under Grant Agreement No. PID2021-127890NB-I00. L.L. acknowledges support by the “Ramón y Cajal” Grant No. RYC2021-031880-I funded by MCIN/AEI/10.13039/501100011033 and the European Union-”NextGenerationEU”. The work of R.R.-G. was partially supported through Grant No. PID2022-136228NB-C22 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE and ”ERDF A way of making Europe.”